OpenStudy (blackops2_143):

What is the value of the expression 1/3^−5 ? 1/15 15 243 1/243

OpenStudy (astrophysics):

Hint: \[\huge x^n = \frac{ 1 }{ x^{-n} }\]

OpenStudy (something):

\[\frac{ 1 }{ \frac{ 1 }{ 3 }^{5} }\]

OpenStudy (something):

the negative becomes the divisor and put it under 1

OpenStudy (blackops2_143):

thx

OpenStudy (something):

no problem